Acetylene a gas used in welding is periodic Ed be the following reaction of calcium carbide CaC2 with water. How much calcium carbide is needed to produce 75.00g of acetylene


1
Expert's answer
2021-03-26T07:51:03-0400
"CaC_2 + 2H_2O = Ca(OH)_2 + C_2H_2"

"n(C_2H_2) = m\/M = 75.0g\/26(g\/mol) = 2.885 mol"

"n(C_2H_2) = n(CaC_2) = 2.885 mol"

"m(CaC_2) = nM = 2.885mol*64g\/mol = 184.64 g"